OpenStudy (anonymous):

Solutions to: h(x)=6x2+4x+3 Fan/medal for help. :)

OpenStudy (anonymous):

6x^2**

OpenStudy (anonymous):

what do u want?

OpenStudy (anonymous):

To find the solutions to the function.

OpenStudy (anonymous):

of the function*

OpenStudy (anonymous):

mean max or min?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

do u mean maximum value of it?

OpenStudy (anonymous):

No. There are two complex solutions to this, and I want to find out what they are.

OpenStudy (anonymous):

ok \[\frac{ -4\pm i2\sqrt{14} }{ 12 }\]

OpenStudy (anonymous):

How'd you get that?

OpenStudy (anonymous):

\[\Delta= b ^{2}-4ac\] \[x=\frac{ -b \pm \sqrt{\Delta} }{ 2a }\]

OpenStudy (ranga):

The solutions for the quadratic equation \[ax ^{2} + bx + c = 0 \] is: \[\frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]

OpenStudy (anonymous):

\[ax ^{2} + bx + c\]

OpenStudy (anonymous):

I know that but that's not what we're learning right now. The lesson is completing the square.

OpenStudy (anonymous):

u mean complex numbers?

OpenStudy (anonymous):

No, the lesson is completing the square. But, in the complex numbers lesson, they said that if the discriminant is negative, there will be two complex solutions to the function.

OpenStudy (anonymous):

The discriminant for h(x) is -56.

OpenStudy (anonymous):

ya right

OpenStudy (anonymous):

what is wrong with it?

OpenStudy (anonymous):

I need to know how you got your answer. I don't understand HOW to find the solutions. If you just tell me, I'm not learning anything.

OpenStudy (ranga):

\[\sqrt{-56} = i \sqrt{4x14} = 2i \sqrt{14}\] where i is defined as \[\sqrt{-1}\]

OpenStudy (anonymous):

ya right

OpenStudy (anonymous):

\[\sqrt{-56}=\sqrt{-4\times14}=i2\times \sqrt{14}\]

OpenStudy (ranga):

4 times 14 not the letter "x" above

OpenStudy (anonymous):

Okay, then what?

OpenStudy (anonymous):

Waaaaaaaait. Where'd you get the 14?

OpenStudy (anonymous):

56 = 4 * 14

OpenStudy (anonymous):

Ohhhhhhhhhh. Okay. But what happens to the rest of the function-?

OpenStudy (anonymous):

u know \[\sqrt{-1}= i\]

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

so \[\sqrt{-56}= \sqrt{-1\times4\times14}= i2\times \sqrt{14}\]

OpenStudy (anonymous):

But where's the 12 come from in your solution? And there is supposed to be two solutions, not just one.

OpenStudy (anonymous):

a=6 2a=2*6 = 12

OpenStudy (anonymous):

there is two solutions mention the signs in the formula

OpenStudy (anonymous):

So the other solution is negative rather than positive?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

What does it look like? I promise it's my last question.

OpenStudy (anonymous):

ok \[x1 = \frac{ -4 + i2\sqrt{14} }{ 12}\] \[x2 =\frac{ -4 - i2\sqrt{14} }{ 12 }\]

OpenStudy (anonymous):

got??

OpenStudy (anonymous):

Yup. Thank you. :)

OpenStudy (anonymous):

well